OpenStudy (anonymous):

The brilliant Indian mathematician Ramanujan immediately answered a a question from the English mathematician Hardy. Hardy mentioned the number 1729 which was the number of a cab he had just travelled in ..He said it was'nt a very interesting number but the Indian instantly said it was and defined its unique property. Any suggestions about what it might be? It involves perfect cubes.

OpenStudy (turingtest):

Oh, I read this once :) I think its the lowest number that can be written as the sum of something... about perfect cubes

OpenStudy (anonymous):

Hardy's Words "I remember once going to see him when he was ill at Putney. I had ridden in taxi cab number 1729 and remarked that the number seemed to me rather a dull one, and that I hoped it was not an unfavorable omen. "No," he replied, "it is a very interesting number; it is the smallest number expressible as the sum of two cubes in two different ways." Source: Wikipedia

OpenStudy (turingtest):

Aww, prove it!

OpenStudy (anonymous):

Smallest positive number, he probably meant.

OpenStudy (anonymous):

yes - that guy was phenominal with numbers - and Hardy was probably the best english mathematician ecer

OpenStudy (anonymous):

His later work with Laurel was also quite funny.

OpenStudy (anonymous):

Hmm what about Euler?

OpenStudy (turingtest):

I heard a lot of Ramanujans work was wrong, though he was definitely great.

OpenStudy (anonymous):

Euler wasn't english.

OpenStudy (anonymous):

Oh Sorry

OpenStudy (anonymous):

1^3 + 12^3 = 1729 9^3 + 10^3 = 1729

OpenStudy (turingtest):

proof it's the smallest?

OpenStudy (anonymous):

lol ktklown

OpenStudy (anonymous):

that would be a good exercise turing - i suppose you could prove it using proof by exhaustion

OpenStudy (turingtest):

You mean an list of 1728 numbers that don't work?

OpenStudy (anonymous):

Ramanajan didn't lie proof - completely opposite to hardy's approach to math - that s why they made such a good leam

OpenStudy (turingtest):

I know, I heard Ramanujan got divinely inspired from Krishna as to the sum of some series, which Hardy hated because he had to go and prove it, lol! Plus Hardy was so anti-Platonism. What's that called, formalism? Ramanujan's approach is more about feeling the answer, which makes little sense from a formalist perspective, right?

OpenStudy (anonymous):

yep - but amazingly he was correct most of the time. Unfortunately, he died in his early thirties.

OpenStudy (anonymous):

This problem is wrong but given right numbers one way to approach is to represent any point on the line as a function of sine and cosine. consider @ as theta Given A(x1,y1) and slope tan@ you can find a point B on the line r units away from A as Bx = x1+rcos@ By=y1+rsin@. Such representation will help in this case because we have relation between points on the line with varying distances.

OpenStudy (turingtest):

Care to elaborate on that method for calculating the smallest number that can be written as the sum of two cubes two different ways? I don't get it.

OpenStudy (anonymous):

I am sorry the explanation was meant for different question I accidently posted it here.

OpenStudy (turingtest):

Oh, I was gonna say...

OpenStudy (anonymous):

for this a^3+b^3=c^3+d^3. give it to super computer to get first five sets of a,b,c,d integer values

OpenStudy (turingtest):

...or give it to Ramanujan :D

OpenStudy (anonymous):

true such a genius only once in a century

Parth (parthkohli):

1729 is called the Hardy-Ramanujan number which can be expressed in the form of the sum of two perfect cubes in two different ways.

Parth (parthkohli):

\[12^3 + 1^3 = 9^3 + 10^3\]I remember how there are other numbers like this too.